In a certain card game, you are dealt two cards face up. You then bet on whether a third card is dealt between the other two cards. What is the probability of winning your bet if you are dealt:
A-a 5 and 7?
B-a pair of 9's
C-a 5 and a queen

Answer by stanbon(75887)

In a certain card game, you are dealt two cards face up. You then bet on whether a third card is dealt between the other two cards. What is the probability of winning your bet if you are dealt:
A-a 5 and 7?
There are 4 6's that are winning cards
There are 52-6 = 46 non-winning cards
-------------------------------------------
Ans: 4/46
-------------------
B-a pair of 9's
There are zero winning cards
There are 50 non-winning cards
Ans: 0/50 = 0
----------------------------
C-a 5 and a queen
There are 6*4 = 24 winning cards
There are 26 non-winning cards
Ans: 24/26
=======================
Cheers,
Stan H.